Dan invested 20,000 dollars and earned 900 dollars in income in the first year. How much income would he have earned if he had invested 25,000 dollars in the same account?
Answer by rfer(16322)

20000/900=25000/x
20000x=22500000
x=22500000/20000
x=$1125